a car makes at a displacement of 100m toward east and then 200m towards north.find tne magnitude and direction of the resultant?
 

Grade:

4 Answers

Shivam
9 years ago
R = [(200)2 + (100)2]1/2    in the dirn of N-E
Shivam
9 years ago
sorry dirn is 100i+200j/[50000]1/2
Ankit Jaiswal
8 years ago
resultant displacement is (100+2002)1/2 =(10000+40000)1/2=500001/2=100(5)½
the direction is North east
 
ali
6 years ago
223.7m tan-1 [2]  n of e …..................................................................//////////////////////////////////////k